OpenStudy (cutiecomittee123):

How many solutions does the equation cos(6x)=1/2 have on an interval of (1,2pi]? I just need a few pointers on this one.

OpenStudy (rizags):

First, try pretending the 6x is a y, then get some primitive solutions and go from there

OpenStudy (cutiecomittee123):

Okay I need more info because I know that cos is y on the unit circle when trying to find solutions. Is this basically just saying "How many places does cos=1/2 on the unit cirlce?

OpenStudy (rizags):

yes, and what are those places.

OpenStudy (cutiecomittee123):

but how is the 6x tied into the answer? like what does it mean in terms of how many places cos=1/2

OpenStudy (cutiecomittee123):

there are only 2 places where cos is equal to 1/2 on the unit circle

OpenStudy (rizags):

Which are: \[\frac{5\pi}{3}, \frac{\pi}{3}\]

OpenStudy (cutiecomittee123):

Yeah

OpenStudy (rizags):

So from there, divide both by 6, like so:\[\frac{\pi}{3}\times\frac{1}{6}=\frac[\pi}{18}\]

OpenStudy (rizags):

Sorry that should say: \[\frac{5\pi}{18}, \frac{\pi}{18}\]

OpenStudy (rizags):

both of those, when plugged in for x, will give you 0.5

OpenStudy (rizags):

However, your interval was (1,2pi] correct?

OpenStudy (cutiecomittee123):

yes

OpenStudy (cutiecomittee123):

@rizags

OpenStudy (lochana):

still confusing?

OpenStudy (lochana):

general solutions for y = cosx is\[x = 2n\pi \pm sin^{-1}y\]

OpenStudy (lochana):

in your case it is\[6x = 2n\pi \pm cos^{-1}(1/2)\]

OpenStudy (lochana):

\[x = \frac{2n\pi \pm \pi/3}{6}\]

OpenStudy (lochana):

now find solutions in (1,pi] sub set by putting n=0,1,2....

OpenStudy (lochana):

those will be solutions

OpenStudy (cutiecomittee123):

So there imore than 2 solutions?

OpenStudy (cutiecomittee123):

I am a little confused on your last comment.

OpenStudy (lochana):

I didn't calculate answers. so I can't say how many solutions will be there. you can substitute n with 0,1,2... and see answers are in the given set.

OpenStudy (cutiecomittee123):

There was 12 solutions in all